OpenStudy (anonymous):

mike wants to make meat loaf. His recipe uses a total of 8 pounds of meat. If he uses a 3 to 1 ratio of beef to pork,how much pork will he use? enter your answer as a mixed number in simplest terms

OpenStudy (dangerousjesse):

\[8\div4=x\]

OpenStudy (dangerousjesse):

Solve for x first.

OpenStudy (dangerousjesse):

Then you will solve: \[8-x*3\]

OpenStudy (dangerousjesse):

\[x*3=n\]

OpenStudy (dangerousjesse):

\[8-x*3=d\]

OpenStudy (dangerousjesse):

\[\frac{ n }{ d }\] will be your answer.